OpenStudy (anonymous):

Identify the sequence graphed below and the average rate of change from n = 1 to n = 3.

OpenStudy (anonymous):

(1, 20) (3, 5). I already know the rate of change is -15/2

OpenStudy (anonymous):

I need help picking which option it is, though

OpenStudy (anonymous):

an = 20(1/2)n - 1; average rate of change is -15/2 or an = 10(1/2)n - 1; average rate of change is -15/2

jimthompson5910 (jim_thompson5910):

the point (1,20) represents the term when n = 1, then a1 = 20 ie the first term is 20

jimthompson5910 (jim_thompson5910):

(3,5) means "the third term is 5"

OpenStudy (anonymous):

OOOH okay! so the answer is the one which starts with a 20, right??

jimthompson5910 (jim_thompson5910):

to find the average rate of change from n = 1 to n = 3, you just find the slope of the line through those two points